OpenStudy (anonymous):

evaluate the expression Arctan(-sqrt 3)

OpenStudy (jdoe0001):

plug it in your calculator :)

OpenStudy (anonymous):

i know but i had a question before that and the question was evaluate the expression arctan(-sqrt 3) what is the differences between the As?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

between the A's?

OpenStudy (jdoe0001):

\(\bf tan^{-1}(\sqrt{3})=\theta\implies tan(\theta)=\sqrt{3}\) if you check your calculator, it has a \(\bf [tan^{-1}]\) button